OpenStudy (anonymous):

-> Need Help <- How do I solve this problem? 3(3x+6)=4(7x-1)

OpenStudy (phi):

As a first step, I would "distribute" the 3 on the left side... that means multiply every term inside the parens by 3 also distribute the 4 on the right side what do you get ?

OpenStudy (anonymous):

9x+18 = 28x-4

OpenStudy (phi):

yes, now add -9x to both sides, like this -9x + 9x + 18 = 28x - 9x -4 can you simplify that ?

OpenStudy (phi):

on the left side you have 9 x's take away 9 x's on the right side you have 28 x's take away 9 x's

OpenStudy (anonymous):

18=19-4

OpenStudy (phi):

not just 19, it is 19x (28 x's take away 9 x's leaves 19 x's) 18 = 19x -4 now add +4 to both sides and simplify what do you get ?

OpenStudy (phi):

like this 18+4= 19x -4 + 4